# Nuanced Safety in Generative AI: How Demographics Shape Responsiveness to Severity

### v1

@@ 6 Conclusion
We formulated non-parametric metrics to assess raters’ responsiveness to the severity of content violations, addressing limitations in existing approaches. Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to severity when assessing AI-generated content. These findings underscore the value of our metrics for understanding and improving the accuracy and reliability of generative AI content safety evaluations. This provides a foundation for a nuanced understanding of raters’ preferences and the expression of those preferences.
7 Limitations and Future Work
In this work, we based our analysis on a simplified data model that assumes the existence of some true underlying severity $V$. We used binary labels obtained from experts as a reference for this true underlying exceeding some threshold. So, a straightforward extension of our work will be to explore the paradigm where we obtain binary labels from crowd raters themselves as a reference for whether the individual-specific perception of severity exceeds some threshold. Doing so will allow us to more directly compare the responsiveness of raters to severity as captured by the collective judgement of the diverse crowd and enable disentangling of rater-specific biases in scale usage from differences in perceptions of severity.
Note that severity is a multi-faceted and intricate concept, the perception of which may be unique to every individual. In future work, we will explore the more complex data model with rater-specific perceptions of safety. Larger datasets will allow to estimate the parameters of our data model (similar to e.g. Homan et al. (2023)), enabling deeper analyses and inferences of rater behaviours and severity perceptions with our proposed metrics.
Impact Statement
This work raises important ethical considerations regarding the potential reliance on human raters from diverse demographic backgrounds, particularly under-represented ones, for evaluating harmful content. Repeated exposure to such content can cause significant emotional distress and trauma (Steiger et al., 2021). To mitigate this risk, we advocate for the deployment of active learning strategies (Kirk et al., 2022) to identify the most informative items and reduce the quantity of harmful content raters need to evaluate. Additionally, this work contributes to making simulated raters (Thomas et al., 2025) more reliable and better aligned with human raters, which can also help alleviate the burden on human raters, particularly for tasks involving high volumes of potentially harmful content.
Appendix A Comparison of Metrics via Simulation
@@
We compare the behaviour of our proposed metrics, monotonic precision area and weighted recall area, with that of traditional metrics like Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve, and area under the ROC curve by simulating different scoring patterns. For the simulations, we assume that the $\mathcal{F}^{\prime}$ in our data model specified in 3.1 is linear, i.e., $R_{ij}=V_{i}+b_{j}$. We further assume that severities $V$ and biases $b$ are both normally distributed. We have 30 crowd raters in our simulations who use a 0 (not harmful) to 4 (completely harmful) Likert scale to score 1000 items. We consider three different scoring patterns:
@@
    Downward shift, where the crowd raters systematically shift a proportion of their scores in the range 2 to 4 downwards.
@@
Figure 6 presents the distribution of crowd rater scores from the three scoring patterns. We compute 7 metrics for the three scoring patterns: monotonic precision area, weighted recall area, their harmonic mean, Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve, and area under the ROC curve. In order to obtain binary labels $U$ for computing the metrics, we simulate 30 experts with varying thresholds $T$. Each expert is allocated a percentile $p$ randomly drawn from a normal distribution with range $[50,90]$; the expert gives a binary score of 1 to any item with $V$ in the top $p$ percentile and 0 otherwise. As before, we obtain the binary labels $U$ for the items by assigning the individual binary scores of every expert to each item via replication. This simulation setup is very general and does not impose any other constraints on observed data.
@@
Figure 6: Distribution of scores from three different scoring patterns of crowd raters in our simulations.
Figure 7: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve (aucpr), and area under the ROC curve (auroc) for three different scoring patterns of simulated crowd rater population. All confidence intervals are within ±$\pm 0.01$.
Figure 7 shows the average metric values for the three different patterns. We see that monotonic precision area does not differ hugely between the normal scoring pattern and the pattern with systematic downward shift. This is expected since relative ordering is largely undisturbed by a systematic downward shift in scores. However, weighted recall area decreases significantly because a systematic downward shift hurts discrimination at each score. Traditional metrics show trends similar to weighted recall area since they focus mostly on the ability to discriminate but do not reflect well the ability to stochastically order. This is further validated when we look at the metrics for the conservative scoring pattern. As expected, weighted recall area and traditional metrics do not differ hugely between the normal scoring pattern and the conservative scoring pattern since high severity items still have a score greater than 0 while others have a score of 0. On the contrary, monotonic precision area drops significantly due to the disruption in stochastic ordering.
Appendix B Group cohesion metrics
Table 1 and Table 2 show the inter-rater agreement among different demographic-based rater groupings. We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024).
@@ Appendix B Group cohesion metrics
Appendix C Detailed Result Tables and Other Plots
This section provides detailed numerical values for the plots shown in the paper.
Table 3: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for trisectional demographic groups of crowd raters when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
Group MPA WRA HM
White GenX Man 0.3533 0.5312 0.4244
South-Asian GenX Man 0.3108 0.5762 0.4038
Black GenZ Woman 0.3049 0.5694 0.3971
White GenX Woman 0.2762 0.5655 0.3711
White GenZ Woman 0.2844 0.5299 0.3701
East-Asian Millennial Man 0.2591 0.5368 0.3495
East-Asian GenZ Woman 0.2518 0.5468 0.3448
South-Asian Millennial Woman 0.2487 0.5473 0.3420
White Millennial Man 0.2449 0.5227 0.3335
Latinx Millennial Man 0.2361 0.5395 0.3285
South-Asian Millennial Man 0.2368 0.5319 0.3277
Black Millennial Woman 0.2367 0.5277 0.3268
Black GenX Woman 0.2328 0.5355 0.3245
Latinx GenX Woman 0.2289 0.5367 0.3209
Latinx Millennial Woman 0.2254 0.5414 0.3183
Latinx GenX Man 0.2213 0.5195 0.3104
White GenZ Man 0.2141 0.5416 0.3069
Black Millennial Man 0.2154 0.5258 0.3056
South-Asian GenX Woman 0.2133 0.5274 0.3038
White Millennial Woman 0.2140 0.5041 0.3005
Black GenX Man 0.2136 0.4981 0.2990
South-Asian GenZ Woman 0.2043 0.5265 0.2944
East-Asian Millennial Woman 0.2058 0.4989 0.2914
Black GenZ Man 0.2009 0.5136 0.2888
East-Asian GenX Woman 0.1996 0.5149 0.2877
East-Asian GenZ Man 0.1891 0.4986 0.2742
South-Asian GenZ Man 0.1750 0.5150 0.2612
Latinx GenZ Man 0.1717 0.5064 0.2564
East-Asian GenX Man 0.1614 0.5133 0.2456
Latinx GenZ Woman 0.1209 0.4915 0.1941
Table 4: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the entire crowd rater population on the three violation types when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
Violation type MPA WRA HM
Sexual 0.4572 0.6933 0.5510
Violent 0.3176 0.4878 0.3847
Bias 0.2155 0.4698 0.2955
Table 5: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean for the top-level demographic groups on the three violation types when using experts to obtain binary labels $U$. All confidence intervals are within ±$\pm 0.01$.
@@ Appendix C Detailed Result Tables and Other Plots
Figure 8: $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three different violation types when using experts to obtain binary labels $U$.
Appendix D Example of prompt-image pairs
Figure 9: This image (generated as a response to the query ”stubbed toe nothing gory” received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e. 2) followed by 10 out 25 who think it is mostly unsafe, (i.e. 3). With the calibration provided from demographics groups we can see that these are mostly women raters from a range of ethnicities and age groups
In this section we provide examples from the dataset of images that demonstrate the differences in responsiveness to severity across different demographic groups’ safety evaluation scores, for example figures 9 and 10.
@@ Appendix D Example of prompt-image pairs
Figure 11: This table shows the harmfulness scores provided by raters from different demographic trisections for the image in Figure 9.

### v4

@@ 7 Conclusion
We formulated non-parametric metrics to analyze the nuanced differences in safety feedback of raters expressed via ordinal scales, addressing the limitations in existing approaches. Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to the severity of violations when providing safety feedback on AI-generated content. These findings underscore the value of our metrics in improving the interpretability and quality of feedback used to align AI systems in multi-cultural contexts.
Appendix A Proofs for the two properties
To prove the inequalities for the two properties, let $\mathcal{T}=\{t_{1},\dots,t_{n}\}$ be the set of all thresholds $T$.
A.1 Proof for stochastic ordering property
We have that
$P(U=1|S=s_{1})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{1},T=t)P(T=t)$ (5)
and
$P(U=1|S=s_{2})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{2},T=t)P(T=t)$ (6)
given that $S$ and $T$ are independent by construction. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence
$P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$ (7)
A.2 Proof for discrimination property
For every $t\in\mathcal{T}$, we have that
$P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ (8)
which means, for any given $t$, we have that
$P(S\geq s|U=1,T=t)\geq P(S\geq s|U=0,T=t)$ (9)
Applying Bayes’ rule here, for any given $t$, we get
$\frac{P(S\geq s,U=1,T=t)}{P(U=1,T=t)}\geq\frac{P(S\geq s,U=0,T=t)}{P(U=0,T=t)}$ (10)
Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent by construction, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then,
$P(S\geq s,U=0,T=t)=P(S\geq s)c(t)-a(t)$ (11)
and
$P(U=0,T=t)=c(t)-b(t)$ (12)
So, for any given $t$, the inequality becomes
$\frac{a(t)}{b(t)}\geq\frac{P(S\geq s)c(t)-a(t)}{c(t)-b(t)}$ (13)
Cross-multiplying and summing the inequalities over all $t\in\mathcal{T}$, we have that
$\sum_{t\in\mathcal{T}}a(t)\geq P(S\geq s)\sum_{t\in\mathcal{T}}b(t)$ (14)
This yields $P(S\geq s,U=1)\geq P(S\geq s)P(U=1)$, and hence, $P(S\geq s|U=1)\geq P(S\geq s)$. The same inequality, upon substitutions, also yields $P(S\geq s|U=0)\leq P(S\geq s)$. Therefore,
$P(S\geq s|U=1)\geq P(S\geq s|U=0)$ (15)
Appendix B Metric implementation
Below we provide code implementing our proposed metrics. The code can be run cheaply on CPUs and has a complexity of $\mathcal{O}(NKR)$, where $N$ is the number of items and $R$ is the number of raters.
{python}
import numpy, sklearn
def get_mpa_wra_for_rater( likert_scores: list, binary_reference: list, likert_length: int ) -> tuple[float]: """ Args: - likert_scores: array of Likert scores from a crowd rater - binary_reference: array of binary reference U - likert_length: length of the Likert scale [0, …, K]
Returns - tuple of floats for the crowd rater (MPA, WRA, harmonic mean) """ # convert to numpy arrays reference_arr = numpy.array(binary_reference) scores_arr = numpy.array(likert_scores)
one_counts, zero_counts, counts = [], [], [] for k in range(0, likert_length): scores_flag = scores_arr == k count_at_k = scores_flag.sum() one_count_at_k = (reference_arr * scores_flag).sum() zero_count_at_k = count_at_k - one_count_at_k counts.append(count_at_k) one_counts.append(one_count_at_k) zero_counts.append(zero_count_at_k)
# compute Y_so at Likert scores and then MPA y_so = [] norm = numpy.ceil(likert_length / 2) * numpy.floor(likert_length / 2) for k in range(0, likert_length): demr, numr = numpy.array(counts[:k]), numpy.array(one_counts[:k]) prev_precisions = numpy.maximum.accumulate(numr[demr != 0] / demr[demr != 0]) y = 0 if counts[k] and len(prev_precisions): precision_at_k = one_counts[k] / counts[k] y = (precision_at_k - prev_precisions).sum() y_so.append(y) y_so.append(0.) mpa = sklearn.metrics.auc(range(likert_length + 1), y_so) / norm
# compute Y_d at Likert scores and then WRA y_d = [] U_0, U_1 = (1 - reference_arr).sum(), reference_arr.sum() for k in range(0, likert_length): safe_recall_below_k = sum(zero_counts[:k]) / U_0 if U_0 else 0. unsafe_recall_at_k = one_counts[k] / U_1 if U_1 else 0. y = safe_recall_below_k * unsafe_recall_at_k y_d.append(y) y_d.append(0.) wra = sklearn.metrics.auc(range(likert_length + 1), y_d)
hm = (2 * mpa * wra) / (mpa + wra) return (mpa, wra, hm)
Appendix C Comparison of metrics via simulations
We compare the behaviour of our proposed metrics, monotonic precision area (MPA) and weighted recall area (WRA), against that of Kendall’s $\tau$, Spearman’s $\rho$, area under the PR curve (AUCPR), and area under the ROC curve (AUROC) by simulating different scoring patterns.
C.1 Simulation setup
@@
For the simulations, we assume that $V^{j}_{i}=V_{i}+b_{j}$, where severities $V_{i}$ and rater-specific tendencies $b_{j}$ are both normally distributed. We have 30 crowd raters in our simulations who use a 0 (not harmful) to $K$ (completely harmful) Likert scale to score 1000 items. We consider three different scoring patterns:
@@
    Downward shift, where the crowd raters systematically shift a proportion of their scores in the range 2 to $K$ downwards.
@@
Figure 5 presents the distribution of crowd rater scores from the three scoring patterns for $K$ = 4. We compute 7 metrics in total for the three scoring patterns: MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, Spearman’s $\rho$, AUCPR, and AUROC. In order to obtain binary reference $U$ for computing the metrics, we simulate 30 trained raters with varying strictness. Each trained rater is allocated a percentile $p$ randomly drawn from a truncated normal distribution with range $[50,90]$; the trained rater gives a binary score of 0 to any item with $V$ in the bottom $p$ percentile and 1 otherwise. As before, we obtain the binary reference $U$ for the items by assigning the individual binary scores of every trained rater to each item via replication. This simulation setup is very general and does not impose any other constraints on observed data.
@@
Figure 5: Distribution of scores from three different scoring patterns of crowd raters in our simulations when $K=4$.
C.2 Simulation results
Figure 6 shows the average metric values for the three different patterns. We see that MPA does not differ hugely between the normal scoring pattern and the pattern with systematic downward shift. This is expected since relative ordering is largely undisturbed by a systematic downward shift in scores. However, WRA decreases significantly because a systematic downward shift hurts discrimination at each score. Traditional metrics show trends similar to WRA since they focus on the ability to discriminate but do not reflect well the ability to stochastically order. They only capture ordinal relationships (concordance / discordance) but not cardinal relationships (the magnitude of differences). This is further validated when we look at the metrics for the conservative scoring pattern. As expected, WRA and traditional metrics do not differ hugely between the normal scoring pattern and the conservative scoring pattern since higher severity items still have a score greater than 0 while others have a score of 0. On the contrary, MPA drops significantly due to the disruption in stochastic ordering.
Figure 6: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman’s $\rho$, area under the pr curve (aucpr), and area under the roc curve (auroc) for the three different scoring patterns of simulated crowd rater population. All confidence intervals are within ±$\pm 0.01$.
C.3 Robustness verification
We further repeated the simulations with scales of longer lengths, i.e., $K\in[6,24]$ and found that the trends observed above remain the same even for scales of longer lengths.
Appendix D Dataset characteristics
The distribution of scores provided by crowd raters in the dataset are shown in figure 7.
Tables 1 and 2 show the inter-rater agreement among different demographic-based rater groupings. We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024).
@@ Appendix D Dataset characteristics
Appendix E Top-level demographic groups on violation types
We compare the scoring patterns of top-level demographic groups on each violation type. Table 3 gives monotonic precision area (MPA), weighted recall area (WRA), and their harmonic mean (HM) for top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference. Looking at the ethnic groups, we see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations. For bias violations, Latinx and Black ethnicities have the highest responsiveness, which is understandable given that the guidelines of expert raters may be geared towards the experiences of these groups when it comes to bias and stereotypes. Looking at the gender groups, we do not see significant differences in responsiveness to severity across the three violation types. Finally, looking at the age groups, we see that the oldest age group, GenX, exhibits the highest MPA for sexual and bias violations. This suggests that GenX raters are the most responsive to granular variations in severity $V^{g}$ of these violations as captured by the guidelines of expert raters.
Table 3: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
@@ Appendix E Top-level demographic groups on violation types
Appendix F Examples of prompt-image pairs
Figures 8 and 9 are examples of prompt-image pairs from the dataset that demonstrate the differences in responsiveness to severity across different demographic groups’ safety feedback.
Figure 8: This image (generated as a response to the query "stubbed toe nothing gory" received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e. 2) followed by 10 out 25 who think it is mostly unsafe, (i.e. 3). When looking at the top-level demographic groups, we see that it is women raters who provide the most unsafe feedback with higher scores.
@@ Appendix F Examples of prompt-image pairs
Figure 10: This table shows the feedback provided by raters from different demographic trisections for the image in Figure 8.
Appendix G Recommendations and visualizations for observed trends
In table 4, we provide some recommendations on steps to take for specific trends that may be observed in monotonic precision area (MPA) and weighted recall area (WRA) of individual crowd raters or rater groups when the binary reference $U$ is either obtained from trained raters (capturing $V^{g}$) or crowd raters themselves (capturing $V^{c}$).
Possible reason(s) Recommended steps
Both MPA and WRA are low There is significant misalignment between the $V^{j}$ of the rater or the rater group and $V^{g}$ or $V^{c}$. The rater or the rater group has misunderstood the ordinal scale, or the binary reference $U$ itself is significantly noisy. Audit the scores of the rater or the rater group and the binary reference $U$ itself by inspecting them on a small set of golden items. Consider revising the instructions given to the raters, both crowd raters as well as trained raters.
[…]
Table 4: Recommended steps to take for various observed trends in monotonic precision area (mpa) and weighted recall area (wra).
Furthermore, in figure 11, we visualize some distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both MPA and WRA of rater $j$ are low, (b) MPA of rater $j$ is low, and (c) WRA of rater $j$ is low. In the case where both MPA and WRA are low, we can see that the perception of rater $j$ is largely orthogonal to the reference. In the case where MPA is low, we can see that patterns like extreme responses can significantly hurt stochastic ordering. And in the case where WRA is low, we can see that a tendency towards the central score can significantly hurt discrimination. We note that these distributions are not exhaustive and are only meant to be illustrative.
Figure 11: Some possible distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both mpa and wra of rater $j$ are low, (b) mpa of rater $j$ is low, and (c) wra of rater $j$ is low. Red dots represent items with $U=1$ and blue dots represent items with $U=0$. Here, the rater uses a 0 to 4 Likert scale, and the dotted horizontal lines demarcate the regions where score $s\in\{0,1,2,3,4\}$ is the most probable.
Appendix H Validation on another real-world dataset
To further validate our metrics on more real-world data, we experiment with the toxicity dataset of Kumar et al. (2021). The authors compiled a dataset of 107,620 comments rated for toxicity by diverse raters of different demographic backgrounds with varying religious and political leaning. Each rater provided a score on a 0-4 ordinal scale where 0 means not toxic and 4 means extremely toxic. An aggregate binary label for each comment was computed by taking the median of the rater scores and comparing if it is above the midpoint score of 2 or not. The authors noted in their analysis that religious leaning, age, and sexuality had the most statistically significant impact on the raters’ scores out of all the demographic axes considered.
In figure 12, we show monotonic precision area (MPA), weighted recall area (WRA), the harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional groups of raters based on their religious leaning, their sexuality, and their age group. As in the original paper, the binary reference $U$ is obtained by taking the median, excluding the rater group being evaluated, and comparing it to the midpoint score of 2. Once again, we note that Kendall’s $\tau$ and AUROC continue to track WRA closely as both of them capture aspects of concordance. But again, they do not reflect well the ability to stochastically order as MPA does.
Figure 12: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional groups of diverse raters. All confidence intervals are within ±$\pm 0.01$. The metric curves are vertically-aligned to start at 0 for ease of comparison. Legend gives the values by which the curves are translated.
Particularly, we observe that trisections where religion is important consistently have lower MPA ($p$-value < 0.05 under permutation test), and consequently, lower HMs as well. This indicates that higher scores from these trisections correspond the least to higher severity $V^{c}$ as captured by the collective judgment of the crowd. This is further corroborated by the fact that Kumar et al. (2021) observed in their own analysis that religious leaning, even when mild, significantly increased the odds of higher toxicity scores from raters. The same trends hold for trisections where sexuality is LGBTQ+, both according to our metrics and according to the analysis of Kumar et al. Lastly, we see that trisections with age group 18-24 consistently have the highest HMs ($p$-value < 0.05), indicating that they are the most responsive to both granular variations and broader distinctions in $V^{c}$. The potential reasoning for this can again be found in the analysis of Kumar et al. (2021) who noted that younger participants may be the most familiar with presence of slangs or certain style of attacks in the comments since the comments were sampled from websites on which such participants happen to be the most active age group.